The equation x^2+(k+2)x+2k=0 has two distinct real roots.
Find the possible values of k.

Answer by ikleyn(52747)

.
HINT: the discriminant must be positive.


Refering to general quadratic equation form  ax^2 + bx + c = 0,  the discriminant is d = b^2-4ac.


In your case b = (k+2), c = 2k,  a = 1.


So, the discriminant is d = b^2 - 4ac = (k+2)^2 - 4*1*2k = k^2 + 4k + 4 - 8k = k^2 - 4k + 4 = (k-2)^2.


The requirement d > 0 means that k =/= 2.


It is your answer:  k =/= 2.

Solved.

Notice that if k = 2, then the two real roots merge to one.


--------------
On discriminant and the roots of a quadratic equation see the lessons
    - Introduction into Quadratic Equations
    - PROOF of quadratic formula by completing the square
in this site.


Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the topic "Quadratic equations".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.